OpenStudy (anonymous):

How do you do this type of problem?: x+2y=0

myininaya (myininaya):

what are you trying to do to it?

OpenStudy (anonymous):

Solve it.

OpenStudy (anonymous):

And then I graph it. But I know how to graph.

OpenStudy (anonymous):

\[x+2y=0\]

myininaya (myininaya):

solve it for what? do you want x-intercepts, y-intercepts. finding these will be useful in graphing.

OpenStudy (anonymous):

Ok I need to find M and B if that makes sense. So yeah x and y intercepts.

myininaya (myininaya):

you need to write in slope intercept form: y=mx+b solve the equation above for y by subtracting x first let me know what you have after you do this

OpenStudy (anonymous):

0+2y=-1?

myininaya (myininaya):

what happen to the x?

myininaya (myininaya):

2y=-x

OpenStudy (anonymous):

and thats it?

myininaya (myininaya):

no its not in this form y=mx+b yet what do we need to do next

OpenStudy (anonymous):

Divide by two?

myininaya (myininaya):

right so we have y=-x/2 so what is m=? and what is b=?

OpenStudy (anonymous):

m=1\1 b=0

OpenStudy (anonymous):

Is that right?

myininaya (myininaya):

m=-1/2 and b=0

OpenStudy (anonymous):

ok. How would you do y=-4

myininaya (myininaya):

y=-4 is a horizontal line going through -4 on the y axis

OpenStudy (anonymous):

Ok so let me give you a problem. Me work it out and will you tell me if its right?

OpenStudy (anonymous):

My next problem is: 2x+3y+6

OpenStudy (anonymous):

y=2\3=2 m=2\3 b= 2

myininaya (myininaya):

what did you write for the y=mx+b form

OpenStudy (anonymous):

y=2\3x=2

myininaya (myininaya):

2x+3y=6 -2x -2x 3y=-2x+6 _ _ _ 3 3 3 y=-2x/3+2 so the slope is m=-2/3 the y intercept is b=2

OpenStudy (anonymous):

One more equation I just dont get.

OpenStudy (anonymous):

x-2y=-4